# Multiplicity of powers of path ideals of a line graph

### v1

@@ Abstract.
Let $S=K[x_{1},\ldots,x_{n}]$ and let $I$ be the $t$-path ideal of the line graph $L_{n}$ with $n$-vertices. It is shown that the set of associated prime ideals of $I^{s}$ is equal to the set of minimal prime ideals of $I$ for all $s\geq 1$, and we provide an explicit description of these prime ideals. Additionally, as the main contribution of this paper, we derive an explicit formula for the multiplicity of $S/I^{s}$ for all $s\geq 1$, revealing that it is a polynomial in $s$ from the beginning.
@@ 1. Introduction
The edge ideal was introduced and studied by Villarreal in [18]. Let $G$ be a finite simple graph with vertex set $V=\{x_{1},\cdots,x_{n}\}$ and edge set $E$. The ideal generated by all quadratic monomials $x_{i}x_{j}$ such that $\{x_{i},x_{j}\}\in E$ is called the edge ideal of $G$, denoted by $I(G)$. Since then, the study of edge ideals of simple graphs, particularly the properties of their powers, has become a vibrant and active topic in combinatorial commutative algebra. For more detailed information, refer to the exposition paper [2] and the references therein.
In 1999, Conca and De Negri introduced the concept of a $t$-path ideal in their work [7], which serves as a generalization of the definition of an edge ideal. For an integer $2\leq t\leq n$, a $t$-path of the graph $G$ is a sequence of $(t-1)$ distinct edges $\{x_{i_{1}},x_{i_{2}}\},\{x_{i_{2}},x_{i_{3}}\},\cdots,$$\{x_{i_{t-1}},x_{i_{t}}\}$ of $G$ such that the vertices $x_{i_{1}},x_{i_{2}},\ldots,x_{i_{t}}$ are pairwise distinct. Such a path is also denoted by $\{x_{i_{1}},\ldots,x_{i_{t}}\}$ for short. The $t$-path ideal $I_{t}(G)$ associated to $G$ is defined as the squarefree monomial ideal
@@
in the polynomial ring $K[x_{1},\ldots,x_{n}]$. In the recent years, some algebraic properties of path ideals have been investigated extensively, see for instance [1, 4, 5, 13, 15, 16].
@@ 1.1. Associated prime ideals
Let $R$ be a Noetherian ring and $M$ an finitely generated non-zero $R$-module. A prime ideal $\operatorname{\mathfrak{p}}$ of $R$ is called an associated prime ideal of $M$ if there exists $a\in M$ such that $\operatorname{\mathfrak{p}}=(0:_{R}a)$. The set of all associated prime ideal of $M$ is denoted by $\mathrm{Ass}_{R}(M)$. It is known that $\mathrm{Ass}_{R}(M)$ is a non-empty and finite set. If $I$ is an ideal of $R$, we denote $\mathrm{Ass}_{R}(R/I)$ by $\mathrm{ass}(I)$. In 1979, Brodmann [6] proved the remarkable result that there exists an integer $s_{0}$ such that $\mathrm{ass}(I^{s})=\mathrm{ass}(I^{s_{0}})$ for all $s\geq s_{0}$. This result serves as a good example of the concept in commutative algebra that ideals exhibit favorable behavior asymptotically. (Another example illustrating this idea is provided later in the paper regarding the multiplicity of an ideal.) Naturally, Brodmann’s theorem also inspires several questions, such as determining the value of $s_{0}$ and determine the prime ideals in $\mathrm{ass}^{\infty}(I)$? Here, $\mathrm{ass}^{\infty}(I):=\mathrm{ass}(I^{s_{0}})$. We refer to [8] and the references therein for a comprehensive introduction to this topic. If $s_{0}=1$, then $I$ is said to be normally torsion-free. More formally, we give the following definition following [10].
@@
An ideal $I$ of a Noetherian ring $R$ is said to be normally torsion-free if $\mathrm{ass}(I^{s})=\mathrm{ass}(I)$ for all $s\geq 1$.
We will show that both $I_{t}(L_{n})$ and $I_{t}(L_{n})^{\vee}$ are normally torsion-free and determine the prime ideals in $\mathrm{ass}^{\infty}(I_{t}(L_{n}))$.
@@ 1.2. Multiplicity
Let $S=K[x_{1},\ldots,x_{n}]$ be a polynomial ring that standard graded over a field $K$ and let $M$ be a finitely generated graded $S$-module. The Hilbert function $H_{M}$ is defined as the function
@@
where $M_{k}$ is the degree $k$ component of $M$. The Hilbert series $\operatorname{HS}(M,z)$ is defined as
$\operatorname{HS}(M,z):=\sum_{i\in\mathbb{Z}}\dim_{\mathbb{K}}M_{i}z^{i}.$
If we assume that $\dim M=d$, it is a celebrated result of David Hilbert that $\operatorname{HS}(M,z)$ is a rational function of the following form
$\operatorname{HS}(M,z)=\frac{\mathbf{Q}(M,z)}{(1-z)^{d}}.$
Here, $\mathbf{Q}(M,z)\in\mathbb{Q}[z^{-1},z]$ is a Laurent polynomial such that $\mathbf{Q}(M,1)\neq 0$. As a consequence of this, the Hilbert function $H_{M}$ is of polynomial type, meaning that there exists a polynomial $p_{M}\in\mathbb{Q}[x]$ such that $H_{M}(k)=p_{M}(k)$ for all $k\gg 0$. The polynomial $p_{M}$ is referred to as the Hilbert polynomial of $M$.
@@
The integer coefficients $e_{i}(M)$ are called the Hilbert coefficients of $M$. The first Hilbert coefficient $e_{0}(M)$ holds a special significance, it is also known as the multiplicity of $M$ and is often denoted by $\operatorname{mult}(M)$.
Let $I$ be a graded ideal of $S$ such that $\dim M/I^{s}M=d$. It was shown in [12, Theorem 1.1] that $e_{i}(M/I^{s}M)$ is of polynomial type in $s$ of degree $\leq n-d+i$ for $i=1,2,\ldots,d$. Recall that a function $f:\mathbb{N}\rightarrow\mathbb{Q}$ is of polynomial type of degree $d$ if there exists a polynomial $p\in\mathbb{Q}[z]$ of degree $d$ such that $f(n)=p(n)$ for all $n\gg 0$. To the best of our knowledge, there are currently no known non-trivial examples of graded ideals for which the multiplicity of their powers has been explicitly computed. In this paper, we will provide an explicit formula for $\operatorname{mult}(S/I_{t}(L_{n})^{s})$.
@@ 1.3. Structure of this paper
Let $I$ denote the $t$-path ideal of the line graph $L_{n}$. This paper is organized as follows. In Section 2 we demonstrate that both $I$ and its Alexander dual $I^{\vee}$ are normally torsion-free and provide a complete characterization of all the ideals in $\mathrm{ass}^{\infty}(I)$. Additionally, we establish that $S/I$ is sequentially Cohen-Macaulay and determine its projective dimension. In the final section, which constitutes the main contribution of this paper, we present an explicit formula for the multiplicity of $S/I^{s}$ for all $s\geq 1$, revealing that it is a polynomial in $s$ from the beginning.
1.4. Conventions
In the paper, we always assume $n,t$ are integers with $n\geq t\geq 1$. The polynomial ring $K[x_{1},\ldots,x_{n}]$ will be denoted by $S$. We denote $L_{n}$ as the line graph of length $n-1$ and $I_{t}(L_{n})$ as the $t$-path ideal of $L_{n}$. Let $G(I_{t}(L_{n}))$ denote the minimal monomial generating set of $I_{t}(L_{n})$. Specially, $G(I_{t}(L_{n}))=\{u_{1},\ldots,u_{n-t+1}\}$, where $u_{i}=x_{i}x_{i+1}\cdots x_{i+t-1}$ for $i=1,\ldots,n-t+1$. For a positive integer $a$, the set $\{1,\ldots a\}$ is denoted by $[a]$. Finally, for a finite set $A$, $|A|$ refers to the number of $A$.
@@ 2. Associated primes of $I_{t}(L_{n})^{s}$
Let $I$ denote the $t$-path ideal $I_{t}(L_{n})$ of the line graph $L_{n}$. In this section, we show that both $I$ and its Alexander dual $I^{\vee}$ are normally torsion-free and provide an explicit description of minimal prime ideals of $I$. Furthermore, We show that $S/I$ is sequentially Cohen-Macaulay and determine its projective dimension.
@@
and the Stanley Reisner $I_{\Delta}$ ideal of $\Delta$ is defined as
@@
We may look upon $I_{t}(L_{n})$ as the facet ideal of some simplicial complex. For this, we put
$F_{i}:=\{i,\ldots,i+t-1\}\subseteq[n]\mbox{ for }i=1,\ldots,n-t+1$
and let $\Delta_{n,t}$ denote the simplicial complex $\langle F_{1},\ldots,F_{n-t+1}\rangle$. It is immediate to see that the path ideal $I_{t}(L_{n})$ is the facet ideal $I(\Delta_{n,t})$ of $\Delta_{n,t}.$ The following lemma reveals an interesting property of $\Delta_{n,t}$, which we will not use in our subsequent discussion.
Recall a simplicial complex $\Delta$ is called a flag complex if its Stanley-Reisner ideals is generated by monomials of degree 2, or equivalently, all minimal non-faces of $\Delta$ have exactly two elements. Here, a non-face of $\Delta$ refers to a subset of $[n]$ which does not belong to $\Delta$.
@@
Conversely, let $\{i_{1},\ldots,i_{k}\}$ be a non-face of $\Delta_{n,t}$ with $i_{1}<\cdots<i_{k}$. Then if $i_{k}-i_{1}<t$, then $\{i_{1},\ldots,i_{k}\}\subseteq F_{i_{1}}$, a contradiction. This implies $i_{k}-i_{1}\geq t$ and so $x_{F}\in I$. This proves $I\supseteq I_{\Delta_{n,t}}.$
Recall that a facet $F$ of a simplicial complex $\Delta$ is called a leaf if either $F$ is the only facet of $\Delta$, or there exists a facet $G\neq F$ such that $H\cap F\subseteq G\cap F$ for any facet $H\neq F$. A leaf $F$ of $\Delta$ is said to be good if the sets $F\cap H$, where $H\in\mathcal{F}(\Delta)$ is linearly ordered. A simplicial complex $\Delta$ is a simplicial forest if every nonempty subcollection of $\Delta$ has a leaf. A connected simplicial forest is called a simplicial tree. Here, a subcollection of $\Delta$ refers to a simplicial complex whose set of facets is a subset of $\mathcal{F}(\Delta).$ The following observation forms the foundation of all the results in this section.
@@
For all $0<t\leq n$, $\Delta_{n,t}$ is a simplicial tree.
@@
We now present the first main result of this section.
@@
It is shown in [14, Theorem 2.2 and Theorem 2.5] that if a simplicial complex $\Delta$ has no special odd cycles, then the facet ideal $I(\Delta)$ and its Alexander dual $I(\Delta^{*})=I(\Delta)^{\vee}$ are normally torsion-free. On the other hand, by [14, Theorem 3.2], if $\Delta$ is a simplicial forest, then $\Delta$ has no special cycles. Now, the result follows by Lemma 2.2.
By applying [10, Theorem 1.4.6], we immediately obtain the following corollary.
@@
Let $\leq t\leq n$. Then the Rees algebra $\mathcal{R}(I_{t}(L_{n}))$ is a normal Cohen-Macauly domain.
Proof.
By [14, Corollary 1.6].
In the following, we will determine all the minimal prime ideals of $I_{t}(L_{n})$, which will be also useful in the next section. To do this, we introduce some other notation. If $F\subseteq[n]$, we denote by $\operatorname{\mathfrak{p}}_{F}$ the prime ideal generated by $x_{i}$ with $i\in F$.
@@
Let $C_{n,t}$ denote the set of subsets $\{i_{1},\ldots,i_{r}\}$ of $[n]$ with $1\leq i_{1}<\cdots<i_{r}\leq n$ and the following conditions are satisfied:
@@
Recall a vertex cover of the simplicial complex $\Delta$ is a subset $C\subset[n]$ such that $C\cap F\neq\emptyset$ for all $F\in{\mathcal{F}}(\Delta)$. The vertex cover $C$ is called a minimal vertex cover, if no proper subset of $C$ is a vertex cover. It is well-known that $\operatorname{\mathfrak{p}}$ is a minimal prime ideal of $I(\Delta)$ if and only if $\operatorname{\mathfrak{p}}=\operatorname{\mathfrak{p}}_{C}$ for some minimal vertex cover of $\Delta$.
Theorem 2.7.
@@
$\{\operatorname{\mathfrak{p}}_{F}\colon\;F\in C_{n,t}\}.$
@@
It is equivalent to showing that $C_{n,t}$ consists of all minimal vertex cover of $\Delta_{n,t}$. Let $F\in C_{n,t}$. We first show that $F$ is a vertex cover of $\Delta_{n,t}$. Given a facet $F_{i}$ of $\Delta_{n,t}$, where $11\leq i\leq n-t+$ 1. If $i\in F$, then $i\in F\cap F_{i}$ and so $F\cap F_{i}\neq\emptyset$. Suppose that $i\notin F$. If $i<i_{1}$ then $i_{1}\in F_{i}\cap F$ and so $F_{i}\cap F\neq\emptyset$. If $i>i_{1}$, then, since $i_{r}\geq n-t+1$, there exists $11\leq j\leq r-$ 1 such that $i_{j}<i<i_{j+1}$. Since $i_{j+1}-i<i_{j+1}-i_{j}\leq t$, it follows that $i_{j+1}\in F_{i}$ and thus $F\cap F_{i}\neq\emptyset$. This proves $F$ is a vertex cover of $\Delta_{n,t}$. Next, we show that $F$ is minimal, that is, $F\setminus\{i_{j}\}$ is not a vertex cover of $\Delta_{n,t}$ for $j=1,\ldots,r$. But, this is clear by noticing that $F_{1}\cap(F\setminus\{i_{1}\})=\emptyset$, $F_{i_{j-1}+1}\cap(F\setminus\{i_{j}\})=\emptyset$ for $j=2,\ldots,r-1$ and $F_{n-t+1}\cap(F\setminus\{i_{r}\})=\emptyset$. Therefore, $F$ is a minimal vertex cover of $\Delta_{n,t}$ indeed.
@@
Let $2\leq t\leq n$. Then $\mathrm{ass}(I_{t}(L_{n})^{s})=\{\operatorname{\mathfrak{p}}_{F}\colon\;F\in C_{n,t}\}$ for all $s\geq 1$.
@@
The Alexander dual $I_{t}(L_{n})^{\vee}$ of $I_{t}(L_{n})$ has the minimal generating set $G=\{X_{F}\colon\;F\in C_{n,t}\}$. Assume $n=(t+1)p+q$ with $0\leq q\leq t$. Then the largest degree of minimal generators of $I_{t}(L_{n})^{\vee}$ is given by
@@
Denote $\operatorname{Deg}(I_{t}(L_{n})^{\vee})$ as $D$. Consider the following infinite sequence
@@
Let $a_{m}$ denote the $m$-th number of this sequence. Namely, $a_{1}=1,a_{2}=t+1,a_{3}=t+2$ and so on. Thus, $a_{2k}=kt+k$ and $a_{2k+1}=kt+k+1$ for $k\geq 1$. Given $F=\{i_{1},\ldots,i_{r}\}\in C_{n,t}$ with $1\leq i_{1}<\cdots<i_{r}\leq n$, we claim that

### v2

@@ Abstract.
Let $S=\mathbb{K}[x_{1},\ldots,x_{n}]$ and let $I$ be the $t$-path ideal of the line graph $L_{n}$ with $n$-vertices. It is shown that the set of associated prime ideals of $I^{s}$ is equal to the set of minimal prime ideals of $I$ for all $s\geq 1$, and we provide an explicit description of these prime ideals. Additionally, as the main contribution of this paper, we derive an explicit formula for the multiplicity of $S/I^{s}$ for all $s\geq 1$, revealing that it is a polynomial in $s$ from the beginning.
@@ 1. Introduction
Let $G$ be a finite simple graph with vertex set $V=\{x_{1},\ldots,x_{n}\}$ and edge set $E$. The ideal generated by all quadratic monomials $x_{i}x_{j}$ such that $\{x_{i},x_{j}\}\in E$ is called the edge ideal of $G$, denoted by $I(G)$. The edge ideal was introduced and studied by Villarreal in [21]. Since then, the study of edge ideals of simple graphs, particularly the properties of their powers, has become a vibrant topic in combinatorial commutative algebra. For more detailed information in this direction, we refer to the exposition paper [3] and the references therein.
In 1999, Conca and De Negri introduced in their work [9] the concept of a $t$-path ideal, which serves as a generalization of an edge ideal. For an integer $2\leq t\leq n$, a $t$-path of the graph $G$ is a sequence of $(t-1)$ distinct edges $\{x_{i_{1}},x_{i_{2}}\},\{x_{i_{2}},x_{i_{3}}\},\cdots,$$\{x_{i_{t-1}},x_{i_{t}}\}$ of $G$ such that the vertices $x_{i_{1}},x_{i_{2}},\ldots,x_{i_{t}}$ are pairwise distinct. Such a path is also denoted by $\{x_{i_{1}},\ldots,x_{i_{t}}\}$ for short. The $t$-path ideal $I_{t}(G)$ associated to $G$ is defined as the square-free monomial ideal
@@
in the polynomial ring $\mathbb{K}[x_{1},\ldots,x_{n}]$. In the recent years, some algebraic properties of path ideals have been investigated extensively, see for instance [1, 5, 6, 15, 17, 18].
@@ 1.1. Associated prime ideals
Let $S$ be a Noetherian ring and $M$ a finitely generated non-zero $S$-module. A prime ideal $\operatorname{\mathfrak{p}}$ of $S$ is called an associated prime ideal of $M$ if there exists $x\in M$ such that $\operatorname{\mathfrak{p}}=(0:_{S}x)$. The set of all associated prime ideal of $M$ is denoted by $\mathrm{Ass}_{S}(M)$. It is known that $\mathrm{Ass}_{S}(M)$ is a non-empty finite set. If $I$ is an ideal of $S$, we denote $\mathrm{Ass}_{S}(S/I)$ by $\mathrm{ass}(I)$. In 1979, Brodmann [7] proved the remarkable result that there exists an integer $s_{0}$ such that $\mathrm{ass}(I^{s})=\mathrm{ass}(I^{s_{0}})$ for all $s\geq s_{0}$. This result serves as a good example of the concept in commutative algebra that ideals exhibit favorable behavior asymptotically. Naturally, Brodmann’s theorem also inspires several questions, such as determining the value of $s_{0}$ and determine the prime ideals in $\mathrm{ass}^{\infty}(I):=\mathrm{ass}(I^{s_{0}})$. We refer to [10] and the references therein for a comprehensive introduction to this topic. If $s_{0}=1$, then $I$ is said to be normally torsion-free. Formally, we give the following definition following [12].
@@
An ideal $I$ of a Noetherian ring $S$ is said to be normally torsion-free if $\mathrm{ass}(I^{s})=\mathrm{ass}(I)$ for all $s\geq 1$.
We will show that both $I_{t}(L_{n})$ and $I_{t}(L_{n})^{\vee}$ are normally torsion-free and determine the prime ideals in $\mathrm{ass}(I_{t}(L_{n}))$.
@@ 1.2. Multiplicity
Let $S=\mathbb{K}[x_{1},\ldots,x_{n}]$ be a polynomial ring over a field $\mathbb{K}$ that is standardly graded and let $M$ be a finitely generated graded $S$-module. The Hilbert function $H_{M}$ of $M$ is defined as the function
@@
where $M_{k}$ is the degree $k$ component of $M$. The Hilbert series $\operatorname{HS}(M,z)$ is defined to be
$\operatorname{HS}(M,z):=\sum_{k\in\mathbb{Z}}H_{M}(k)z^{k}.$
If we assume $\dim M=d+1$, it is a celebrated result of David Hilbert that $\operatorname{HS}(M,z)$ is a rational function of the following form
$\operatorname{HS}(M,z)=\frac{Q(M,z)}{(1-z)^{d+1}}.$
Here, $Q(M,z)\in\mathbb{Q}[z^{-1},z]$ is a Laurent polynomial such that $Q(M,1)\neq 0$. As a consequence of this result, the Hilbert function $H_{M}$ is of polynomial type of degree $d$, meaning that there exists a polynomial $p_{M}(z)\in\mathbb{Q}[z]$ of degree $d$ such that $H_{M}(k)=p_{M}(k)$ for all $k\gg 0$. The polynomial $p_{M}(z)$ is referred to as the Hilbert polynomial of $M$.
@@
The integer coefficients $e_{i}(M)$ for $i=0,\ldots,d$ are called the Hilbert coefficients of $M$. Among these, the first Hilbert coefficient $e_{0}(M)$ holds particular significance. It is also known as the multiplicity of $M$ and and is frequently denoted by $\operatorname{mult}(M)$.
Recall that a function $f:\mathbb{N}\rightarrow\mathbb{Q}$ is of polynomial type of degree $d$ if there exists a polynomial $p(z)\in\mathbb{Q}[z]$ of degree $d$ such that $f(k)=p(k)$ for all $k\gg 0$. Let $I$ be a graded ideal of $S$ such that $\dim M/IM=d$. It was shown in [14, Theorem 1.1] that $e_{i}(M/I^{s}M)$ is of polynomial type in $s$ of degree $\leq n-d+i$ for $i=1,2,\ldots,d$. To the best of our knowledge, there are currently no known non-trivial examples of graded ideals for which the multiplicity of their powers has been explicitly computed. In this paper, we will provide an explicit formula for the multiplicity of $S/I_{t}(L_{n})^{s}$.
@@ 1.3. Structure of this paper
In the paper, we always assume that $n,t$ are integers with $n\geq t\geq 1$. The polynomial ring $\mathbb{K}[x_{1},\ldots,x_{n}]$ will be denoted by $S$. We use $L_{n}$ to denote the line graph with $n$ vertices and let $I_{t}(L_{n})$ be the $t$-path ideal of the line graph $L_{n}$. This paper is organized as follows. In Section 2 we demonstrate that both $I_{t}(L_{n})$ and its Alexander dual $I_{t}(L_{n})^{\vee}$ are normally torsion-free and provide a complete characterization of all the associated prime ideals of $I_{t}(L_{n})$. The final section contains the main contribution of this paper, in which we present an explicit formula for the multiplicity of $S/I_{t}(L_{n})^{s}$ for all $s\geq 1$, revealing that it is a polynomial in $s$ from the beginning.
@@ 2. Associated primes of $I_{t}(L_{n})^{s}$
According to the definition of a $t$-path ideal, the ideal $I_{t}(L_{n})$ has a unique minimal generating set $u_{1},\ldots,u_{n-t+1}$, where $u_{i}=x_{i}\cdots x_{i+t-1}$ for $i=1,\ldots,n-t+1$. In this section, we demonstrate that both $I_{t}(L_{n})$ and its Alexander dual $I_{t}(L_{n})^{\vee}$ are normally torsion-free and offer an explicit description of minimal prime ideals of $I_{t}(L_{n})$. Furthermore, utilizing a novel method, we compute the projective dimension of $S/I_{t}(L_{n})$.
@@
and the Stanley-Reisner $I_{\Delta}$ ideal of $\Delta$ is defined as
@@
Put $F_{i}:=\{i,\ldots,i+t-1\}\subseteq[n]$ for $i=1,\ldots,n-t+1$ and define $\Delta_{n,t}$ to be the simplicial complex generated by $F_{1},\ldots,F_{n-t+1}$. It is evident that the path ideal $I_{t}(L_{n})$ is equal to the facet ideal $I(\Delta_{n,t})$ of $\Delta_{n,t}.$ A simplicial complex $\Delta$ is called a flag complex if its Stanley-Reisner ideals is generated by monomials of degree 2, or equivalently, all minimal non-faces of $\Delta$ have exactly two elements. Here, a non-face of $\Delta$ refers to a subset of $[n]$ which does not belong to $\Delta$. The lemma below uncovers an intriguing characteristic of $\Delta_{n,t}$, although we won’t explore it further in this paper.
@@
Conversely, let $F=\{i_{1},\ldots,i_{k}\}$ be a non-face of $\Delta_{n,t}$ with $i_{1}<\cdots<i_{k}$. Then if $i_{k}-i_{1}<t$, then $\{i_{1},\ldots,i_{k}\}\subseteq F_{i_{1}}$, a contradiction. This implies $i_{k}-i_{1}\geq t$ and so $x_{F}\in I$. This proves $I\supseteq I_{\Delta_{n,t}}.$
A facet $F$ of a simplicial complex $\Delta$ is called a leaf if either $F$ is the only facet of $\Delta$, or there exists a facet $G\neq F$ such that $H\cap F\subseteq G\cap F$ for any facet $H\neq F$. A leaf $F$ of $\Delta$ is said to be good if the sets $F\cap H$, where $H\in\mathcal{F}(\Delta)$ is linearly ordered. A simplicial complex $\Delta$ is a simplicial forest if every nonempty subcollection of $\Delta$ has a leaf. A connected simplicial forest is called a simplicial tree. Here, a subcollection of $\Delta$ refers to a simplicial complex whose set of facets is a subset of $\mathcal{F}(\Delta).$ The following observation forms the foundation of all the results in this section.
@@
For all $0<t\leq n$, the simplicial complex $\Delta_{n,t}$ is a simplicial tree.
@@
We now present the main result of this section.
@@
It was shown in [16, Theorem 2.2 and Theorem 2.5] that if a simplicial complex $\Delta$ has no special odd cycles, then both the facet ideal $I(\Delta)$ and its Alexander dual $I(\Delta^{*})=I(\Delta)^{\vee}$ are normally torsion-free. On the other hand, by [16, Theorem 3.2], if $\Delta$ is a simplicial forest, then $\Delta$ has no special cycles. In view of these facts, the result follows immediately from Lemma 2.2.
We note that the result that $I_{t}(L_{n})$ is normally torsion-free has been established in [19, Proposition 3.2] utilizing a completely different method.
By applying [12, Theorem 1.4.6] and [16, Corollary 1.6] respectively, we immediately obtain the following corollaries.
@@
Let $1\leq t\leq n$. Then the Rees algebra $\mathcal{R}(I_{t}(L_{n}))$ is a normal Cohen-Macauly domain.
In the following, we aim to determine all the minimal prime ideals of $I_{t}(L_{n})$, a task that will also prove useful in the subsequent section. To do this, we introduce some additional notation. Specifically, if $F\subseteq[n]$, we denote by $\operatorname{\mathfrak{p}}_{F}$ the prime ideal generated by $x_{i}$ with $i\in F$. For convenience, we also adopt the following notation.
@@
Let $\mathbf{C}_{n,t}$ be the set of subsets $\{i_{1},\ldots,i_{r}\}$ of $[n]$ with $1\leq i_{1}<\cdots<i_{r}\leq n$ and such that the following conditions are satisfied:
@@
Recall that a vertex cover of the simplicial complex $\Delta$ is a subset $C\subset[n]$ such that $C\cap F\neq\emptyset$ for all $F\in{\mathcal{F}}(\Delta)$. A vertex cover $C$ is called a minimal vertex cover, if no proper subset of $C$ is a vertex cover. It is well-known that $\operatorname{\mathfrak{p}}$ is a minimal prime ideal of $I(\Delta)$ if and only if $\operatorname{\mathfrak{p}}=\operatorname{\mathfrak{p}}_{C}$ for some minimal vertex cover of $\Delta$.
Proposition 2.7.
@@
$\{\operatorname{\mathfrak{p}}_{F}\colon\;F\in\mathbf{C}_{n,t}\}.$
@@
It is equivalent to showing that $\mathbf{C}_{n,t}$ consists of all minimal vertex covers of $\Delta_{n,t}$. Let $F\in\mathbf{C}_{n,t}$. We first show that $F$ is a vertex cover of $\Delta_{n,t}$. Given a facet $F_{i}$ of $\Delta_{n,t}$, where $11\leq i\leq n-t+$ 1. If $i\in F$, then $i\in F\cap F_{i}$ and so $F\cap F_{i}\neq\emptyset$. Suppose that $i\notin F$. If $i<i_{1}$ then $i_{1}\in F_{i}\cap F$ and so $F_{i}\cap F\neq\emptyset$. If $i>i_{1}$, then, since $i_{r}\geq n-t+1$, there exists $11\leq j\leq r-$ 1 such that $i_{j}<i<i_{j+1}$. Since $i_{j+1}-i<i_{j+1}-i_{j}\leq t$, it follows that $i_{j+1}\in F_{i}$ and thus $F\cap F_{i}\neq\emptyset$. This proves $F$ is a vertex cover of $\Delta_{n,t}$. Next, we show that $F$ is minimal, that is, $F\setminus\{i_{j}\}$ is not a vertex cover of $\Delta_{n,t}$ for $j=1,\ldots,r$. But, this is clear by noticing that $F_{1}\cap(F\setminus\{i_{1}\})=\emptyset$, $F_{i_{j-1}+1}\cap(F\setminus\{i_{j}\})=\emptyset$ for $j=2,\ldots,r-1$ and $F_{n-t+1}\cap(F\setminus\{i_{r}\})=\emptyset$. Therefore, $F$ is a minimal vertex cover of $\Delta_{n,t}$ indeed.
@@
This result can be checked using [2, Theorem 3.1]. Combining Theorem 2.3 with Proposition 2.7, we obtain the following result.
@@
Let $2\leq t\leq n$. Then $\mathrm{ass}(I_{t}(L_{n})^{s})=\{\operatorname{\mathfrak{p}}_{F}\colon\;F\in\mathbf{C}_{n,t}\}$ for all $s\geq 1$.
Given a monomial ideal $I$, we use $\operatorname{Deg}(I)$ to denote the maximal degree of minimal generators of $I$. We now determine $\operatorname{Deg}(I_{t}(L_{n})^{\vee})$ for all $1\leq t\leq n$. Recall for a square-free monomial ideal $I=I(\Delta)$, the Alexander dual of $I$ is defined to be the monomial ideal
$I^{\vee}=(x_{C}\colon\;C\mbox{ is a vertex cover of }\Delta).$
@@
The Alexander dual $I_{t}(L_{n})^{\vee}$ of $I_{t}(L_{n})$ has the minimal generating set $G=\{x_{F}\colon\;F\in\mathbf{C}_{n,t}\}$. Assume $n=(t+1)p+q$ with $0\leq q\leq t$. Then the maximal degree of minimal generators of $I_{t}(L_{n})^{\vee}$ is given by
@@
The first statement is immediate from the definition of the Alexander dual. To prove the second statement, we denote $\operatorname{Deg}(I_{t}(L_{n})^{\vee})$ as $D$ for simplicity. Consider the following infinite sequence
@@
Let $a_{m}$ denote the $m$-th number of this sequence. Namely, $a_{1}=1,a_{2}=t+1,a_{3}=t+2$ and so on. Thus, $a_{2k}=kt+k$ and $a_{2k+1}=kt+k+1$ for $k\geq 1$. Given $F=\{i_{1},\ldots,i_{r}\}\in\mathbf{C}_{n,t}$ with $1\leq i_{1}<\cdots<i_{r}\leq n$, we claim that